OpenStudy (anonymous):

What is the sum of a 6-term geometric sequence if the first term is 11, the last term is –11,264 and the common ratio is –4?

OpenStudy (anonymous):

sum=a(1-r)^n/(1-r) where a is the first term r is the common ratio and n is the number of terms

OpenStudy (anonymous):

did i do something wrong ? Sn= a(1 - r^n) ------------- 1- r Sn= 11(1 - (-4)^6) ------------- 1- (-4) Sn= 11(1 - (-4096)) ------------- 5 Sn= 11(4097) ------------- 5 Sn= 11(4097) ------------- 5

ganeshie8 (ganeshie8):

hint : (-4)^6 = 4^6 = 4096

ganeshie8 (ganeshie8):

Sn= a(1 - r^n) ------------- 1- r Sn= 11(1 - (-4)^6) ------------- 1- (-4) Sn= 11(1 - (4096)) ------------- 5 Sn= 11(-4095) ------------- 5

OpenStudy (anonymous):

99099 ?

ganeshie8 (ganeshie8):

im getting -9009

OpenStudy (anonymous):

ohh noo . i see my mistake . your righht . What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250? how do i answer this with out the common ration ?

ganeshie8 (ganeshie8):

we need to find r first

ganeshie8 (ganeshie8):

\(a_8 = 10 \times r^7 = 781250\) => \( r^7 = 78125\) \(r = 5\)

ganeshie8 (ganeshie8):

Now try to find the sum using the usual sum formula

OpenStudy (anonymous):

one sec .

ganeshie8 (ganeshie8):

yo

OpenStudy (anonymous):

im sorry . are you here @ganeshie8 ?

ganeshie8 (ganeshie8):

yea

ganeshie8 (ganeshie8):

its okay :)

OpenStudy (anonymous):

Sn= a(1 - r^n) ------------- 1- 5 Sn= 10(1 - 5^8) ------------- -4 Sn= 10(1 - 390625) ------------- -4 Sn= 10(- 390624) ------------- -4 am i right so far right ?

ganeshie8 (ganeshie8):

yes. keep going

OpenStudy (anonymous):

976560 ?

ganeshie8 (ganeshie8):

correct ! gw :)

OpenStudy (anonymous):

Thankyouu (: . ii need help on few more of these questions. 5. Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August?

ganeshie8 (ganeshie8):

March : 2 times April : 2^2 times May : 2^3 times June : 2^4 times July : 2^5 times Aug : 2^6 times

ganeshie8 (ganeshie8):

So, in August it wud be 5 x 2^6 = ?

OpenStudy (anonymous):

soo 320 ?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

6. A local grocery store stacks the soup cans in such a way that each row has 2 fewer cans than the row below it. If there are 32 cans on the bottom row, how many total cans are on the bottom 14 rows?

ganeshie8 (ganeshie8):

use arithmetic series formula

ganeshie8 (ganeshie8):

first term, a = 32

ganeshie8 (ganeshie8):

common difference, d = -2

ganeshie8 (ganeshie8):

n = 14

ganeshie8 (ganeshie8):

\(Sn = \frac{n}{2}(2a + (n-1)d)\) = ?

OpenStudy (anonymous):

use that equation ? ^^^^^

ganeshie8 (ganeshie8):

yes thats the sum formula for arithmetic series

ganeshie8 (ganeshie8):

give it a try

OpenStudy (anonymous):

Sn=n/2(2a+(n-1)d) Sn=14/2(2(32)+(14-1)-2) Sn=7(64+13-2) Sn=7(75) Sn=525

ganeshie8 (ganeshie8):

Sn=n/2(2a+(n-1)d) Sn=14/2(2(32)+(14-1)(-2)) Sn=7(64+13(-2)) Sn=7(64-26) Sn= ?

OpenStudy (anonymous):

Sn=7(64-26) Sn=7(38) Sn=190

ganeshie8 (ganeshie8):

Sn=7(64-26) Sn=7(38) Sn= ?

OpenStudy (anonymous):

opps i mean 266

ganeshie8 (ganeshie8):

correct !

OpenStudy (anonymous):

thankyou very much (: . i only have 3 more. 8. A fireplace contains 46 bricks along its bottom row. If each row above decreases by 4 bricks, how many bricks are on the 12th row? am i supposed to use the same formula ?

ganeshie8 (ganeshie8):

its arithmetic sequence but this time the question is not about finding sum

ganeshie8 (ganeshie8):

we need to find the 12th term

ganeshie8 (ganeshie8):

\(a = 46\) \(d = -4\) \(a_{12} = ?\)

ganeshie8 (ganeshie8):

use the nth term of arithmetic sequence formula : \(a_n = a + (n-1)d\)

OpenStudy (anonymous):

so it is suppose 46n=46+(n-1)-4d to be

ganeshie8 (ganeshie8):

\(a_{12} = 46 + (12-1)(-4)\) = ?

OpenStudy (anonymous):

a12=46+(12-1)(-4) a12=46+(11)(-4) a12=46+-44 a12=2

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

yesss (:. ok last two ! 7. A major US city reports a 12% increase in decoration sales during the yearly holiday season. If decoration sales were 8 million in 1998, how much did the city report in total decoration sales by the end of 2004?

ganeshie8 (ganeshie8):

So, 1998 to 2004, thats how many years ?

OpenStudy (anonymous):

6

ganeshie8 (ganeshie8):

good, n = 6

ganeshie8 (ganeshie8):

a = 8 million

ganeshie8 (ganeshie8):

12% increase => r = 1.12

OpenStudy (anonymous):

so use the geometric sequence ?

ganeshie8 (ganeshie8):

sales in 2004 = \(\large ar^{6-1}\) = \(\large 8 (1.12)^{5}\) = ?

ganeshie8 (ganeshie8):

^ yes

OpenStudy (anonymous):

ok i got Sn= a(1 - r^n) ------------- 1- r Sn= 8,000,000(1 - 1.12^6) ------------- 1- 1.12 Sn= 8,000,000(1 - 1.97) ------------- 1- 1.12 Sn= 8,000,000(-0.97) ------------- 1- 1.12 Sn= -7760000 ------------- 0.12

ganeshie8 (ganeshie8):

i think the question is oly asking about sales in 2004, so we dont have to find the sum of sales in all years

ganeshie8 (ganeshie8):

just find the sales in 2004, that wud be enough i guess

ganeshie8 (ganeshie8):

Since you found the Sum in all years already, lets finish it.

ganeshie8 (ganeshie8):

Sn= a(1 - r^n) ------------- 1- r Sn= 8,000,000(1 - 1.12^6) ------------- 1- 1.12 Sn= 8,000,000(1 - 1.97) ------------- 1- 1.12 Sn= 8,000,000(-0.97) ------------- 1- 1.12 Sn= -7760000 ------------- -0.12 = ?

OpenStudy (anonymous):

Sn= -7760000 ------------- 0.12 -6466666 I ALSO DID the other formula Sales in 2004 ,,,, ar^6-1 8(1.12)^5 (8)1.76 14

OpenStudy (anonymous):

was that right ? ^^^ last one . Using complete sentences, explain the difference between an exponential function and a geometric series.

ganeshie8 (ganeshie8):

both are right. see ur options and tick which ever exists

ganeshie8 (ganeshie8):

Sn= -7760000 ------------- -0.12 6466666

ganeshie8 (ganeshie8):

its a positive value

OpenStudy (anonymous):

opps . i have to be careful with those signs.

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

can you help with the last one ?

ganeshie8 (ganeshie8):

sure

OpenStudy (anonymous):

Using complete sentences, explain the difference between an exponential function and a geometric series.

ganeshie8 (ganeshie8):

interesting q

OpenStudy (anonymous):

all i know is geometric series is a formula.

ganeshie8 (ganeshie8):

Both exponential function and a geometric sequence increase exponentially. The only difference between them is that a geometric sequence is discrete, but an exponential function is defined everywhere. In other words, a geometric sequence is defined only for few values only, where as an exponential function is defined is smooth and defined everywhere.

ganeshie8 (ganeshie8):